OpenStudy (anonymous):

NEED HELP! DON'T UNDERSTAND! (The is supposed to be a Greek E) 10 E 2^i i=1 A. 2046 B. 2047 C. 1024 D. 2048 E. 1023

OpenStudy (anonymous):

try a

OpenStudy (anonymous):

Can you explain how to do it?

OpenStudy (kirbykirby):

You can make use of this formula: \[\Large\sum_{i=0}^{n-1}ar^i=a\left( \frac{1-r^n}{1-r}\right) \]

OpenStudy (kirbykirby):

\[ \Large \sum_{i=1}^{10}2^i=\sum_{i=0}^{9}2^{i+1}=\sum_{i=0}^92(2^i)\] so \(a = 2, r=2, n=10\) n = 10 since 9 = n -1 => n = 9+1 = 10